OpenStudy (voltron21):

HELP ME!!!!!! For which value of n does the intermediate value theorem guarantee that f(x)= x^3+ 2x + 1 has a zero in the interval (n/2, (n+1)/2)

OpenStudy (anonymous):

what does the intermediate value theorem say?

OpenStudy (voltron21):

a continuous function does not skip vaues

OpenStudy (anonymous):

Nope, that isn't what it says exactly, though that is somewhat of a corollary I suppose.

OpenStudy (anonymous):

i agree

OpenStudy (voltron21):

clearly im clueless thats what my calculus teacher told me

OpenStudy (anonymous):

It says that if \(f(x)\) is continuous on the closed interval \([a,b]\), then there exists a \(c\) where \(f(a) \leq f(c) \leq f(b)\).

OpenStudy (anonymous):

Now, if we're trying to prove a root exists, we'd want \(f(a) \leq 0\) and \(f(b) \geq 0\)

OpenStudy (anonymous):

Then we'd have a root between \([a,b]\)

OpenStudy (voltron21):

i get that there has to be a positive and a negative value i just dont know how to find it in this case given that the intervals have variables

OpenStudy (anonymous):

You need to solve for \(n\).

OpenStudy (voltron21):

so find f(n/2) and f(n+1/2)?

OpenStudy (anonymous):

\[ \frac{n^3}{2^3}+\frac {2n}{2}+1\leq 0 \]

OpenStudy (anonymous):

That is the equation \(f(n/2)\leq 0\). The other one is \(f((n+1)/2) \geq 0\).

OpenStudy (voltron21):

\[f(\frac{ n }{ 2})\le1\]

OpenStudy (voltron21):

?

OpenStudy (voltron21):

im lost again

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[ \frac{n^3}{2^3}+\frac {2n}{2}+1\leq 0 \] \(\color{blue}{\text{End of Quote}}\) Find solutions for \(n\).

OpenStudy (anonymous):

\[ n^3+8n+8\leq 0 \]

OpenStudy (anonymous):

Actually, the easiest thing to do might be simply to find the roots of the actual polynomial first.

OpenStudy (anonymous):

Find roots of \[ x^3+ 2x + 1 \]

OpenStudy (anonymous):

Then you know: \[ \frac n2 \leq r \leq \frac{n+1}2 \]

OpenStudy (anonymous):

Or rather: \[ n\leq 2r\leq n+1 \]

OpenStudy (voltron21):

the root of \[x^3+ 2x+1 \] is \[\pm1\]

OpenStudy (anonymous):

\(-1\) is a root. \(1\) isn't.

OpenStudy (anonymous):

So you could do: \[ n\leq 2(-1)\leq n+1\implies n\leq -2 \leq n+1 \]You could let \(n=-2.5\)

OpenStudy (anonymous):

And test that out and see if it works.

OpenStudy (voltron21):

it didnt work would i just keep plugging in random numbers for n until it works?

OpenStudy (anonymous):

Does \(n\) have ot be a natural number?

OpenStudy (voltron21):

idk my homework is online and it tells me if the answer is wrong and i plugged in n=-2.5 and it said incorrect

OpenStudy (anonymous):

Oh wait!

OpenStudy (anonymous):

\(-1\) wasn't a root. It was \(-\frac 12\)

OpenStudy (voltron21):

how did u figure that?

OpenStudy (anonymous):

So: \[ n\leq -1\leq n+1 \]

OpenStudy (voltron21):

are the numbers you're choosing random!

OpenStudy (voltron21):

?

OpenStudy (voltron21):

oh because its n/2

OpenStudy (anonymous):

No, I already explained my logic.

OpenStudy (anonymous):

Yes. that is why I multiply the root by \(2\)

OpenStudy (voltron21):

ok so n=-1 thank you so much :)

OpenStudy (voltron21):

so this problem is much easier than it seems